# 360DVD: Controllable Panorama Video Generation with 360-Degree Video Diffusion Model

### v1

@@ Abstract
360-degree panoramic videos recently attract more interest in both studies and applications, courtesy of the heightened immersive experiences they engender. Due to the expensive cost of capturing ${}^{\circ}$ panoramic videos, generating desirable panoramic videos by given prompts is urgently required. Recently, the emerging text-to-video (T2V) diffusion methods demonstrate notable effectiveness in standard video generation. However, due to the significant gap in content and motion patterns between panoramic and standard videos, these methods encounter challenges in yielding satisfactory ${}^{\circ}$ panoramic videos. In this paper, we propose a controllable panorama video generation pipeline named 360-Degree Video Diffusion model (360DVD) for generating panoramic videos based on the given prompts and motion conditions. Concretely, we introduce a lightweight module dubbed 360-Adapter and assisted 360 Enhancement Techniques to transform pre-trained T2V models for ${}^{\circ}$ video generation. We further propose a new panorama dataset named WEB360 consisting of ${}^{\circ}$ video-text pairs for training 360DVD, addressing the absence of captioned panoramic video datasets. Extensive experiments demonstrate the superiority and effectiveness of 360DVD for panorama video generation. The code and dataset will be released soon.
@@ 1 Introduction
With the recent advancements in Virtual Reality (VR) technology, 360-degree panoramic videos have been gaining increasing popularity. This video format which offers audiences an immersive experience, is helpful for various applications, including entertainment, education, and communication. To capture details of the entire scene, $360^{\circ}$ videos are typically recorded using an array of high-resolution fisheye cameras that yields a $360^{\circ}\times 180^{\circ}$ field-of-view (FoV) [1], which is quite costly in both time and resources. Therefore, the generation of $360^{\circ}$ panoramic videos is urgently required for border applications, while panoramic video generation receives little attention in studies to date.
@@
Therefore, we propose a specifically designed method named 360-Degree Video Diffusion (360DVD) for generating panorama videos. We first introduce a plug-and-play module named 360-Adapter to address challenge mentioned above. Our 360-Adapter receives zero values or motion conditions (e.g., optical flow) as input and outputs motion features, which are fed into the frozen denoising U-Net at different levels of the encoder. This transformation is aimed at converting the T2V model into a panoramic video generation without altering the foundational generative capabilities. In addition, we introduce 360 Enhancement Techniques including two mechanisms to enhance continuity at both ends of ERPs from both macro and micro perspectives, and a latitude-aware loss function for encouraging the model to focus more on low-latitude regions. Cooperated with carefully designed techniques, our 360DVD generates text-aligned, coherent, high-quality, ${}^{\circ}$ panorama videos with various styles, as shown in Fig. 1.
@@ 2.1 Text-to-Image Diffusion Model
The Denoising Diffusion Probabilistic Model [18, 41, 9, 10] has proven to be highly successful in generating high-quality images, outperforming previous approaches such as generative adversarial networks (GANs)[12, 61], variational autoencoders (VAEs)[21, 40], and flow-based methods [5]. With text guidance during training, users can generate images based on textual input. Noteworthy examples include GLIDE [29], DALLE-2 [33], Imagen [37]. To address the computational burden of the iterative denoising process, LDM [34] conducts the diffusion process on a compressed latent space rather than the original pixel space. This accomplishment has prompted further exploration in extending customization [36, 15], image guidance [56, 59], precise control [62, 28, 27] and protection [64, 60].
@@ 2.2 Text-to-Video Diffusion Model
Despite significant advancements in Text-to-Image (T2I) generation, Text-to-Video (T2V) generation faces challenges, including the absence of large-scale, high-quality paired text-video datasets, the inherent complexity in modeling temporal consistency, and the resource-intensive nature of training. To address these challenges, many works leverage the knowledge from pre-trained T2I models, and they manage training costs by executing the diffusion process in the latent space. Some methods [51, 31, 58, 16, 52] utilize T2I models in zero-shot or few-shot ways. However, these methods often suffer from suboptimal frame consistency due to insufficient training. To address this limitation, another category of T2V diffusion models typically adopts space-time separable architectures. These models [39, 65, 3, 49, 57] inherit spatial operations from pre-trained T2I models, reducing the complexity of constructing space-time models from scratch. Given that most personalized T2I models are derived from the same base one (e.g. Stable Diffusion [34]), AnimateDiff [17] designs a motion modeling module that trained with a base T2I model and could animate most of derived personalized T2I models once for all. There are also efforts focused on enhancing control in T2V models. Gen-1 [14], MCDiff [6], LaMD [19] and VideoComposer [50] introduce diverse conditions to T2V models. Despite these advancements, the aforementioned methods demand extensive training and lack a plug-and-play nature, making it challenging to apply them to a diverse range of personalized T2I models.
@@ 3.1 Preliminaries
Latent Diffusion Model. Given an input signal $x_{0}$, a diffusion forward process in DDPM [18] is defined as:
$p_{\theta}(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t-1}}x_{t-1},\beta_{t}I),$ (1)
@@
for $t=1,...,T$, where $T$ is the total timestep of the diffusion process. A noise depending on the variance $\beta_{t}$ is gradually added to $x_{t-1}$ to obtain $x_{t}$ at the next timestep and finally reach $x_{T}\in\mathcal{N}(0,I)$. The goal of the diffusion model is to learn to reverse the diffusion process (denoising). Given a random noise $x_{t}$, the model predicts the added noise at the next timestep $x_{t-1}$ until the origin signal $x_{0}$:
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t)),$ (2)
@@
for $t=T,...,1$. We fix the variance $\Sigma_{\theta}(x_{t},t)$ and utilize the diffusion model with parameter $\theta$ to predict the mean of the inverse process $\mu_{\theta}(x_{t},t)$. The model can be simplified as denoising models $\epsilon_{\theta}(x_{t},t)$, which are trained to predict the noise of $x_{t}$ with a noise prediction loss:
@@
$\mathcal{L}=\mathbb{E}_{x_{0},y,\epsilon\sim\mathcal{N}(0,I),t}[\|\epsilon-\epsilon_{\theta}(x_{t},t,\tau_{\theta}(y))\|^{2}_{2}],$ (3)
@@
where $\epsilon$ is the added noise to the input image $x_{0}$, $y$ is the corresponding textual description, $\tau_{\theta}(\cdot)$ is a text encoder mapping the string to a sequence of vectors.
@@
Latent Diffusion Model (LDM) [34] executes the denoising process in the latent space of an autoencoder, namely $\mathcal{E}(\cdot)$ and $\mathcal{D}(\cdot)$, implemented as VQ-GAN [20] or VQ-VAE [44] pre-trained on large image datasets. During the training of the latent diffusion networks, an input image $x_{0}$ is initially mapped to the latent space by the frozen encoder, yielding $z_{0}=\mathcal{E}(x_{0})$. Thus, the training objective can be formulated as follows:
@@
$\mathcal{L}=\mathbb{E}_{\mathcal{E}(x_{0}),y,\epsilon\sim\mathcal{N}(0,I),t}[\|\epsilon-\epsilon_{\theta}(z_{t},t,\tau_{\theta}(y))\|^{2}_{2}].$ (4)
@@
In widely-used LDM Stable Diffusion (SD), which our method is based on, $\epsilon_{\theta}(\cdot)$ is implemented with a modified UNet [35] that incorporates four downsample/upsample blocks and one middle block, resulting in four resolution levels within the networks’ latent space. Each resolution level integrates 2D convolution layers as well as self- and cross-attention mechanisms. Text model $\tau_{\theta}(\cdot)$ is implemented using the CLIP [32] ViT-L/14 text encoder.
@@
AnimateDiff. AnimateDiff inflates base SD by adding temporal-aware structures and learning reasonable motion priors from large-scale video datasets. Since the original SD can only process 4D image data batches, while T2V task takes a 5D video tensor in the shape of $batch\times channels\times frames\times height\times width$ as input. It transforms each 2D convolution and attention layer in the original image model into spatial-only pseudo-3D layers by reshaping the $frame$ axis into the $batch$ axis and allowing the network to process each frame independently. Then, the feature maps are reshaped into $(batch\times height\times width)\times frames\times channels$ before going through the proposed motion module to achieve motion smoothness and content consistency. The motion module is inserted at every resolution level of the U-shaped diffusion network, using vanilla temporal transformers consisting of several self-attention blocks operating along the temporal axis. The training objective of AnimateDiff can be written as:
$\mathcal{L}=\mathbb{E}_{\mathcal{E}(x_{0}^{1:N}),y,\epsilon\sim\mathcal{N}(0,I),t}[\|\epsilon-\epsilon_{\theta}(z_{t}^{1:N},t,\tau_{\theta}(y))\|^{2}_{2}],$ (5)
@@
where $x_{0}^{1:N}$ is the sampled video data, $z_{0}^{1:N}$ is the latent code which $x_{0}^{1:N}$ are encoded into via the pre-trained autoencoder, $z_{t}^{1:N}$ is the latent code obtained by perturbing the initial latent code $z_{0}^{1:N}$ with noise at timestep $t$. During training, the pre-trained weights of the base T2I model are frozen to keep its feature space unchanged.
@@ 3.2 WEB360 Dataset
360 Text Fusion. We find that directly using BLIP [23] to label ERP has drawbacks. On one hand, errors may arise due to the distortion caused by the polarities, leading to misidentifications such as labeling “person” as “dog”. On the other hand, the captions generated by BLIP lack granularity, making them insufficient for providing a detailed description of the current scene. Thus, we propose 360 Text Fusion (360TF) method, as shown in Fig. 2. To deal with the irregular distortion of ERP, we turn to less-distorted perspective images. We first project the original ERP image to four non-overlapping perspective images at 0 degrees longitude, with a FoV of 90. The four images are then fed into BLIP to be captioned. By pre-informing ChatGPT about the task and providing examples, these four captions are collectively input to ChatGPT, which then generates a summary of the scene as our final caption. In comparison to directly using BLIP to label the entire image, our 360TF demonstrates a significant advantage in granularity.
@@
During the training process, we first sample a video $x_{0}^{1:N}$ from the dataset. The video is encoded into latent code $z_{0}^{1:N}$ through pre-trained VAE encoder $\mathcal{E}(\cdot)$ and noised to $z_{t}^{1:N}$. Simultaneously, the corresponding text $y$ for the video is encoded using the text encoder $\tau_{\theta}(\cdot)$ of the CLIP. The video is also input into a motion estimation network to generate corresponding motion conditions $C$, which are then fed into the 360-Adapter $\mathcal{F}_{360}(\cdot)$. Finally, noised latent code $z_{t}^{1:N}$, timestep $t$, text embedding $\tau_{\theta}(y)$, and the feature maps $F_{c}$ generated by 360-Adapter are collectively input into the U-Net $\epsilon(\cdot)$ to predict the noise strength added to the latent code. As we aim to preserve the priors learned by SD and AnimateDiff on large datasets, we freeze their weights during the training process. If we use a simple L2 loss term, the training objective is given as follows:
$\mathcal{L}=\mathbb{E}_{\mathcal{E}(x_{0}^{1:N}),y,\epsilon\sim\mathcal{N}(0,I),t}[\|\epsilon-\epsilon_{\theta}(z_{t}^{1:N},t,\tau_{\theta}(y),F_{c})\|^{2}_{2}].$ (6)
@@
In inference, users have the option to selectively provide text prompts and motion guidance to carry out denoising over a total of $T$ steps. Here, we employ DDIM [41] to accelerate the sampling process. The estimated latent code $\hat{z}_{0}^{1:N}$ is then input into a pre-trained VAE decoder to decode the desired $360^{\circ}$ panoramic videos $\hat{x}_{0}^{1:N}$. Due to constraints such as resolution limitations imposed by existing SD and considerations regarding GPU memory usage, the experimental results presented in this paper showcase a resolution of $512\times 1024$. In practical applications, super-resolution methods [42, 24, 8] can be employed to upscale the generated results to the desired size.
360-Adapter. Our proposed 360-Adapter is simple and lightweight, as shown in Fig. 4. The original condition input has the the same resolution as the video of $H\times W$. Here, we utilize the pixel unshuffle [38] operation to downsample it to $H/8\times W/8$. Following that are four 360-Adapter blocks, we depict only one for simplification in Fig. 4. To maintain consistency with the U-Net architecture, the first three 360-Adapter blocks each include a downsampling block. In each 360-Adapter block, one 2D convolution layer and a residual block (RB) with pseudo-3D convolution layers are utilized to extract the condition feature $F^{k}_{c}$. Finally, multi-scale condition features $F_{c}=\{F^{1}_{c},F^{2}_{c},F^{3}_{c},F^{4}_{c}\}$ are formed. Suppose the intermediate features in the U-Net encoder block is $F_{enc}=\{F^{1}_{enc},F^{2}_{enc},F^{3}_{enc},F^{4}_{enc}\}$. $F_{c}$ is then added with $F_{enc}$ at each scale. In summary, the condition feature extraction and conditioning operation of the 360-Adapter can be defined as the following formulation:
$F_{c}=\mathcal{F}_{360}(C),$ (7)
$\hat{F}^{i}_{enc}=F^{i}_{enc}+F^{i}_{c},i\in\{1,2,3,4\}.$ (8)
In the previous description, we omit some details. Our motion condition $C$ is a 5D tensor, assuming its size is $batch\times channels\times frames\times height\times width$. We first reshape it into a 4D tensor of size $(batch\times frames)\times channels\times height\times width$ to allow it to be fed into the 2D convolution layer and restore it to 5D to go through the RB with pseudo-3D convolution layers. Subsequently, in the RB, we employ a $1\times 3\times 3$ pseudo-3D convolution to extract features in the spatial dimension, followed by a $3\times 1\times 1$ pseudo-3D convolution to model information along the temporal dimension. The resulting features are reshaped back to $(batch\times frames)\times channels\times height\times width$ to add the output of the skip connection. Finally, condition features are reshaped back into a 5D vector of size $batch\times channels\times frames\times height\times width$ to align with the U-Net encoder intermediate features.
@@
$\mathcal{L}=\mathbb{E}_{\mathcal{E}(x_{0}^{1:N}),y,\epsilon\sim\mathcal{N}(0,I),t}[\|w(\epsilon-\epsilon_{\theta}(z_{t}^{1:N},t,\tau_{\theta}(y),F_{c}))\|^{2}_{2}],$ (9)
where $w$ is a weight matrix defined as follows:
$w_{i,j}=\cos{(\frac{2i-h+1}{2h}\pi)},$ (10)
where $i\in\left[0,h\right)$, $j\in\left[0,w\right)$, $h$ and $w$ is the height and width of latent code $z_{t}^{1:N}$. The visualized result of $w$ is shown in Fig. 5, where pixels in low and middle latitudes are given more weight during training.
Figure 5: Left: the visualization of weight matrix $w$, brighter colors indicate values closer to 1, while darker colors suggest values closer to 0. Right: a schematic diagram of the latent rotation mechanism. In each iteration, the far left portion of angle $\theta$ is shifted to the far right.
@@
Latent Rotation Mechanism. Because ERPs can be considered as the unfolding of a spherical surface along a meridian, they are meant to be wraparound consistent, implying that their left and right sides are continuous. However, during the process of video generation, the left and right sides are physically separated. Inspired by PanoDiff [47], we employ a latent rotation mechanism to enhance the macroscopic coherence between the left and right ends of the video. During the inference process, we perform a horizontal rotation at an angle of $\theta$ on $z_{t}^{1:N}$ and motion condition $C$, at each denoising step. As illustrated in Fig. 5, the content on the far left is shifted to the far right, where we use $x_{0}^{1}$ to replace $z_{t}^{1:N}$ for a better visual effect of its continuity. During the training process, we also randomly rotate the training videos along with the motion condition by a random angle as a data augmentation strategy.

### v2

@@ Abstract
Panorama video recently attracts more interest in both study and application, courtesy of its immersive experience. Due to the expensive cost of capturing 360 panoramic videos, generating desirable panorama videos by prompts is urgently required. Lately, the emerging text-to-video (T2V) diffusion methods demonstrate notable effectiveness in standard video generation. However, due to the significant gap in content and motion patterns between panoramic and standard videos, these methods encounter challenges in yielding satisfactory 360 panoramic videos. In this paper, we propose a pipeline named 360-Degree Video Diffusion model (360DVD) for generating 360 panoramic videos based on the given prompts and motion conditions. Specifically, we introduce a lightweight 360-Adapter accompanied by 360 Enhancement Techniques to transform pre-trained T2V models for panorama video generation. We further propose a new panorama dataset named WEB360 consisting of panoramic video-text pairs for training 360DVD, addressing the absence of captioned panoramic video datasets. Extensive experiments demonstrate the superiority and effectiveness of 360DVD for panorama video generation. Our project page is at https://akaneqwq.github.io/360DVD/.
footnotetext: This work was supported by National Natural Science Foundation of China under Grant 62372016. (🖂 Corresponding author: Jian Zhang) footnotetext: The authors would like to thank Miguel Lara (reachable at miguel@deforum.art) and Rahmel Jackson (reachable at rahmelsjackson@gmail.com), for their contributions to remarkable LatentLabs360.
@@ 1 Introduction
With the recent advancements in VR technology, 360-degree panoramic videos have been gaining increasing popularity. This video format which offers audiences an immersive experience, is helpful for various applications, including entertainment, education, and communication. To capture details of the entire scene, $360^{\circ}$ videos are typically recorded using an array of high-resolution fisheye cameras that yields a $360^{\circ}\times 180^{\circ}$ field-of-view (FoV) [1], which is quite costly in both time and resources. Therefore, the generation of $360^{\circ}$ panoramic videos is urgently required for border applications, while panoramic video generation receives little attention in studies to date.
@@
Therefore, we propose a specifically designed method named 360-Degree Video Diffusion (360DVD) for generating panorama videos. We first introduce a plug-and-play module named 360-Adapter to address challenge mentioned above. Our 360-Adapter receives zero values or motion conditions (e.g., optical flow) as input and outputs motion features, which are fed into the frozen denoising U-Net at different levels of the encoder. This transformation is aimed at converting the T2V model into a panoramic video generation without altering the foundational generative capabilities. In addition, we introduce 360 Enhancement Techniques including two mechanisms to enhance continuity at both ends of ERPs from both macro and micro perspectives, and a latitude-aware loss function for encouraging the model to focus more on low-latitude regions. Cooperated with carefully designed techniques, our 360DVD generates text-aligned, coherent, high-quality, 360 panorama videos with various styles, as shown in Fig. 1.
@@ 2.1 Text-to-Image Diffusion Model
The Denoising Diffusion Probabilistic Model [17, 39, 9] has proven to be highly successful in generating high-quality images, outperforming previous approaches such as generative adversarial networks (GANs)[11, 57], variational autoencoders (VAEs)[20, 38], and flow-based methods [5]. With text guidance during training, users can generate images based on textual input. Noteworthy examples include GLIDE [27], DALLE-2 [31], Imagen [35]. To address the computational burden of the iterative denoising process, LDM [32] conducts the diffusion process on a compressed latent space rather than the original pixel space. This accomplishment has prompted further exploration in extending customization [34, 14], image guidance [53, 55], precise control [58, 26, 25] and protection [56].
@@ 2.2 Text-to-Video Diffusion Model
Despite significant advancements in Text-to-Image (T2I) generation, Text-to-Video (T2V) generation faces challenges, including the absence of large-scale, high-quality paired text-video datasets, the inherent complexity in modeling temporal consistency, and the resource-intensive nature of training. To address these challenges, many works leverage the knowledge from pre-trained T2I models, and they manage training costs by executing the diffusion process in the latent space. Some methods [48, 29, 54, 15, 49] utilize T2I models in zero-shot or few-shot ways. However, these methods often suffer from suboptimal frame consistency due to insufficient training. To address this limitation, another category of T2V diffusion models typically adopts space-time separable architectures. These models [37, 60, 3, 46] inherit spatial operations from pre-trained T2I models, reducing the complexity of constructing space-time models from scratch. Given that most personalized T2I models are derived from the same base one (e.g. Stable Diffusion [32]), AnimateDiff [16] designs a motion modeling module that trained with a base T2I model and could animate most of derived personalized T2I models once for all. There are also efforts focused on enhancing control in T2V models. Gen-1 [13], MCDiff [6], LaMD [18] and VideoComposer [47] introduce diverse conditions to T2V models. Despite these advancements, the aforementioned methods demand extensive training and lack a plug-and-play nature, making it challenging to apply them to a diverse range of personalized T2I models.
@@ 3.1 Preliminaries
Latent Diffusion Model. Given an input signal $\mathbf{x}_{0}$, a diffusion forward process in DDPM [17] is defined as:
$p_{\theta}(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I}),$ (1)
@@
for $t=1,\dotsc,T$, where $T$ is the total timestep of the diffusion process. A noise depending on the variance $\beta_{t}$ is gradually added to $\mathbf{x}_{t-1}$ to obtain $\mathbf{x}_{t}$ at the next timestep and finally reach $\mathbf{x}_{T}\in\mathcal{N}(0,\mathbf{I})$. The goal of the diffusion model is to learn to reverse the diffusion process (denoising). Given a random noise $\mathbf{x}_{t}$, the model predicts the added noise at the next timestep $\mathbf{x}_{t-1}$ until the origin signal $\mathbf{x}_{0}$:
$p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};{\boldsymbol{\mu}}_{\theta}(\mathbf{x}_{t},t),{\boldsymbol{\Sigma}}_{\theta}(\mathbf{x}_{t},t)),$ (2)
@@
for $t=T,\dotsc,1$. We fix the variance $\boldsymbol{\Sigma}_{\theta}(\mathbf{x}_{t},t)$ and utilize the diffusion model with parameter $\theta$ to predict the mean of the inverse process ${\boldsymbol{\mu}}_{\theta}(\mathbf{x}_{t},t)$. The model can be simplified as denoising models ${\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{t},t)$, which are trained to predict the noise of $\mathbf{x}_{t}$ with a noise prediction loss:
@@
$\mathcal{L}=\mathbb{E}_{\mathbf{x}_{0},\mathbf{y},{\boldsymbol{\epsilon}}\sim\mathcal{N}(0,\mathbf{I}),t}[\left\|{\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{t},t,\boldsymbol{\tau}_{\theta}(\mathbf{y}))\right\|^{2}_{2}],$ (3)
@@
where ${\boldsymbol{\epsilon}}$ is the added noise to the input image $\mathbf{x}_{0}$, $\mathbf{y}$ is the corresponding textual description, $\boldsymbol{\tau}_{\theta}(\cdot)$ is a text encoder mapping the string to a sequence of vectors.
@@
Latent Diffusion Model (LDM) [32] executes the denoising process in the latent space of an autoencoder, namely $\mathcal{E}(\cdot)$ and $\mathcal{D}(\cdot)$, implemented as VQ-GAN [19] or VQ-VAE [42] pre-trained on large image datasets. During the training of the latent diffusion networks, an input image $\mathbf{x}_{0}$ is initially mapped to the latent space by the frozen encoder, yielding $\mathbf{z}_{0}=\mathcal{E}(\mathbf{x}_{0})$. Thus, the training objective can be formulated as follows:
@@
$\mathcal{L}=\mathbb{E}_{\mathcal{E}(\mathbf{x}_{0}),\mathbf{y},{\boldsymbol{\epsilon}}\sim\mathcal{N}(0,\mathbf{I}),t}[\left\|{\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\theta}(\mathbf{z}_{t},t,\boldsymbol{\tau}_{\theta}(\mathbf{y}))\right\|^{2}_{2}].$ (4)
@@
In widely-used LDM Stable Diffusion (SD), which our method is based on, ${\boldsymbol{\epsilon}}_{\theta}(\cdot)$ is implemented with a modified UNet [33] that incorporates four downsample/upsample blocks and one middle block, resulting in four resolution levels within the networks’ latent space. Each resolution level integrates 2D convolution layers as well as self- and cross-attention mechanisms. Text model $\boldsymbol{\tau}_{\theta}(\cdot)$ is implemented using the CLIP [30] ViT-L/14 text encoder.
@@
AnimateDiff. AnimateDiff inflates base SD by adding temporal-aware structures and learning reasonable motion priors from large-scale video datasets. Since the original SD can only process 4D image data batches, while T2V task takes a 5D video tensor as input. It transforms each 2D convolution and attention layer in the original image model into spatial-only pseudo-3D layers. The motion module is inserted at every resolution level of the U-shaped diffusion network, using vanilla temporal transformers consisting of several self-attention blocks operating along the temporal axis. The training objective of AnimateDiff can be written as:
$\mathcal{L}=\mathbb{E}_{\mathcal{E}(\mathbf{x}_{0}^{1:N}),\mathbf{y},{\boldsymbol{\epsilon}}\sim\mathcal{N}(0,\mathbf{I}),t}[||{\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\theta}(\mathbf{z}_{t}^{1:N},t,\boldsymbol{\tau}_{\theta}(\mathbf{y}))||^{2}_{2}],$ (5)
@@
where $\mathbf{x}_{0}^{1:N}$ is the sampled video data, $\mathbf{z}_{0}^{1:N}$ is the latent code which $\mathbf{x}_{0}^{1:N}$ are encoded into via the pre-trained autoencoder, $\mathbf{z}_{t}^{1:N}$ is the latent code obtained by perturbing the initial latent code $\mathbf{z}_{0}^{1:N}$ with noise at timestep $t$. During training, the pre-trained weights of the base T2I model are frozen to keep its feature space unchanged.
@@ 3.2 WEB360 Dataset
360 Text Fusion. We find that directly using BLIP [22] to label ERP has drawbacks. On one hand, errors may arise due to the distortion caused by the polarities, leading to misidentifications such as labeling “person" as “dog". On the other hand, the captions generated by BLIP lack granularity, making them insufficient for providing a detailed description of the current scene. Thus, we propose 360 Text Fusion (360TF) method, as shown in Fig. 2. To deal with the irregular distortion of ERP, we turn to less-distorted perspective images. We first project the original ERP image to four non-overlapping perspective images at 0 degrees longitude, with a FoV of 90. The four images are then fed into BLIP to be captioned. By pre-informing ChatGPT about the task and providing examples, these four captions are collectively input to ChatGPT, which then generates a summary of the scene as our final caption. In comparison to directly using BLIP to label the entire image, our 360TF demonstrates a significant advantage in granularity.
@@
During the training process, we first sample a video $\mathbf{x}_{0}^{1:N}$ from the dataset. The video is encoded into latent code $\mathbf{z}_{0}^{1:N}$ through pre-trained VAE encoder $\mathcal{E}(\cdot)$ and noised to $\mathbf{z}_{t}^{1:N}$. Simultaneously, the corresponding text $\mathbf{y}$ for the video is encoded using the text encoder $\boldsymbol{\tau}_{\theta}(\cdot)$ of the CLIP. The video is also input into a motion estimation network to generate corresponding motion conditions $\mathbf{c}$, which are then fed into the 360-Adapter $\mathcal{F}_{360}(\cdot)$. Finally, noised latent code $\mathbf{z}_{t}^{1:N}$, timestep $t$, text embedding $\boldsymbol{\tau}_{\theta}(\mathbf{y})$, and the feature maps $\mathbf{f}_{360}$ generated by 360-Adapter are collectively input into the U-Net ${\boldsymbol{\epsilon}}(\cdot)$ to predict the noise strength added to the latent code. As we aim to preserve the priors learned by SD and AnimateDiff on large datasets, we freeze their weights during the training process. If we use a simple L2 loss term, the training objective is given as follows:
$\small\mathcal{L}=\mathbb{E}_{\mathcal{E}(\mathbf{x}_{0}^{1:N}),\mathbf{y},{\boldsymbol{\epsilon}}\sim\mathcal{N}(0,\mathbf{I}),t}[||{\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\theta}(\mathbf{z}_{t}^{1:N},t,\boldsymbol{\tau}_{\theta}(\mathbf{y}),\mathbf{f}_{360})||^{2}_{2}].$ (6)
@@
In inference, users have the option to selectively provide text prompts and motion guidance to carry out denoising over a total of $T$ steps. Here, we employ DDIM [39] to accelerate the sampling process. The estimated latent code $\hat{\mathbf{z}}_{0}^{1:N}$ is then input into a pre-trained VAE decoder to decode the desired $360^{\circ}$ panoramic videos $\hat{\mathbf{x}}_{0}^{1:N}$. Due to constraints such as resolution limitations imposed by existing SD and considerations regarding GPU memory usage, the experimental results presented in this paper showcase a resolution of $512\times 1024$. In practical applications, super-resolution methods [40, 8] can be employed to upscale the generated results to the desired size.
360-Adapter. Our proposed 360-Adapter is simple and lightweight, as shown in Fig. 4. The original condition input has the the same resolution as the video of $H\times W$. Here, we utilize the pixel unshuffle [36] operation to downsample it to $H/8\times W/8$. Following that are four 360-Adapter blocks, we depict only one for simplification in Fig. 4. To maintain consistency with the U-Net architecture, the first three 360-Adapter blocks each include a downsampling block. In each 360-Adapter block, one 2D convolution layer and a residual block (RB) with pseudo-3D convolution layers are utilized to extract the condition feature $\mathbf{f}^{k}_{360}$. Finally, multi-scale condition features $\mathbf{f}_{360}=\{\mathbf{f}^{1}_{360},\mathbf{f}^{2}_{360},\mathbf{f}^{3}_{360},\mathbf{f}^{4}_{360}\}$ are formed. Suppose the intermediate features in the U-Net encoder block is $\mathbf{f}_{enc}=\{\mathbf{f}^{1}_{enc},\mathbf{f}^{2}_{enc},\mathbf{f}^{3}_{enc},\mathbf{f}^{4}_{enc}\}$. $\mathbf{f}_{360}$ is then added with $\mathbf{f}_{enc}$ at each scale. In summary, the condition feature extraction and conditioning operation of the 360-Adapter can be defined as the following formulation:
$\mathbf{f}_{360}=\mathcal{F}_{360}(\mathbf{c}),$ (7)
$\hat{\mathbf{f}}^{i}_{enc}=\mathbf{f}^{i}_{enc}+\mathbf{f}^{i}_{360},i\in\{1,2,3,4\}.$ (8)
In the previous description, we omit some details. Our motion condition $\mathbf{c}$ is a 5D tensor, assuming its size is $batch\times channels\times frames\times height\times width$. We first reshape it into a 4D tensor of size $(batch\times frames)\times channels\times height\times width$ to allow it to be fed into the 2D convolution layer and restore it to 5D to go through the RB with pseudo-3D convolution layers. Subsequently, in the RB, we employ a $1\times 3\times 3$ pseudo-3D convolution to extract features in the spatial dimension, followed by a $3\times 1\times 1$ pseudo-3D convolution to model information along the temporal dimension. The resulting features are reshaped back to $(batch\times frames)\times channels\times height\times width$ to add the output of the skip connection. Finally, condition features are reshaped back into a 5D vector of size $batch\times channels\times frames\times height\times width$ to align with the U-Net encoder intermediate features.
@@
$\mathcal{L}=\mathbb{E}_{\mathcal{E}(\mathbf{x}_{0}^{1:N}),\mathbf{y},{\boldsymbol{\epsilon}}\sim\mathcal{N}(0,\mathbf{I}),t}[||\mathbf{W}\odot({\boldsymbol{\epsilon}}-\hat{{\boldsymbol{\epsilon}}}_{\theta})||^{2}_{2}],$ (9)
where $\hat{{\boldsymbol{\epsilon}}}_{\theta}={\boldsymbol{\epsilon}}_{\theta}(\mathbf{z}_{t}^{1:N},t,\boldsymbol{\tau}_{\theta}(\mathbf{y}),\mathbf{f}_{360})$, and $\mathbf{W}$ is a weight matrix used to perform element-wise product, defined as:
$\mathbf{W}_{i,j}=\cos{(\frac{2i-H/8+1}{H/4}\pi)},$ (10)
where $i\in\left[0,H/8\right)$, $j\in\left[0,W/8\right)$, $H/8$ and $W/8$ is the height and width of latent code $\mathbf{z}_{t}^{1:N}$. The visualized result of $\mathbf{W}$ is shown in Fig. 5, where pixels in low and middle latitudes are given more weight during training.
Figure 5: Left: the visualization of weight matrix $\mathbf{W}$, brighter colors indicate values closer to 1, while darker colors suggest values closer to 0. Right: a schematic diagram of the latent rotation mechanism. In each iteration, the far left portion of angle $\theta$ is shifted to the far right.
@@
Latent Rotation Mechanism. Because ERPs can be considered as the unfolding of a spherical surface along a meridian, they are meant to be wraparound consistent, implying that their left and right sides are continuous. However, during the process of video generation, the left and right sides are physically separated. Inspired by PanoDiff [45], we employ a latent rotation mechanism to enhance the macroscopic coherence between the left and right ends of the video. During the inference process, we perform a horizontal rotation at an angle of $\theta$ on $\mathbf{z}_{t}^{1:N}$ and motion condition $\mathbf{c}$, at each denoising step. As illustrated in Fig. 5, the content on the far left is shifted to the far right, where we use $\mathbf{x}_{0}^{1}$ to replace $\mathbf{z}_{t}^{1:N}$ for a better visual effect of its continuity. During the training process, we also randomly rotate the training videos along with the motion condition by a random angle as a data augmentation strategy.